Circular No. 6571

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970111 K. Hurley, Space Sciences Laboratory, Berkeley, on behalf of the Ulys-

ses Gamma-Ray Burst Team; C. Kouveliotou, G. Fishman, and C. Meegan, on behalf of the Gamma Ray Observatory BATSE Team; and J. van Paradijs, University of Amsterdam and University of Alabama in Huntsville, report: “A refined triangulation of GRB 970111 has been done using almost final Ulysses and BATSE data. The resulting position is an annulus whose half-width is 0^{o}.027, centered at α = 11^{h}50^{m}56^{s}, δ = +33^{o}21′.8 (equinox 2000.0), with central radius 49^{o}.981. Combining this with the refined BeppoSAX error circle by in ’t Zand et al. (IAUC 6569) gives an approximately trapezoidal error box with corners at α, δ = 15^{h}28^{m}24^{s}, +19^{o}34′.2; 15^{h}28^{m}28^{s}, +19^{o}36′.04; 15^{h}28^{m}08^{s}, +19^{o}33′.97; 15^{h}28^{m}20^{s}, +19^{o}39′.25. This error box is consistent with the one reported on IAUC 6545 but has an area approximately seven times smaller. Only minor improvements to this error box are expected.”

T. Galama, University of Amsterdam (UoA); R. Strom, Netherlands Foundation for Research in Astronomy, Dwingeloo, and UoA; J. van Paradijs; L. Hanlon and K. Bennett, European Space Agency/ESTEC; C. Kouveliotou, Universities Space Research Association; G. Fishman and C. Meegan, Marshall Space Flight Center, NASA; J. Heise, SRON, Utrecht; and K. Hurley report: “We have made radio observations at 840 MHz (21′′ × 62′′ beamwidth) with the Westerbork array of the revised BeppoSAX/WFC-IPN error box of GRB 970111 (cf. IAUC 6569 and Hurley et al., above) on Jan. 12 (for 1.5 hr), Jan. 13, 14, 16, 20, and 29 (12 hr each day). We also observed at 5 GHz for 6 hr on each of Feb. 26 and 27. The error box contains one weak unresolved radio source (A) at α = 15^{h}28^{m}17^{s}.7, δ = +19^{o}35′39′′ (equinox 2000.0); two sources lie just outside the error box: B, α = 15^{h}28^{m}12^{s}.8, +19^{o}32′50′′; C, α = 15^{h}28^{m}04^{s}.9, δ = +19^{o}35′00′′. The 840-MHz and 5-GHz fluxes in mJy are as follows: source A, 0.6, < 0.5; B, 1.0, < 0.5; C, < 0.5, 0.8 (upper limits are 3σ).”

COMET 81P/WILD 2 Total visual magnitude estimates: Feb. 2.15 UT, 10.0 (R. Keen, Mt.

Thorodin, CO, 0.32-m reflector); 7.05, 9.8 (J. Bortle, Stormville, NY, 0.41m reflector); 14.06, 9.7 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 27.94, 9.9 (A. Pereira, Cabo da Roca, Portugal, 0.25-m reflector); Mar. 1.07, 9.9 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector).

1997 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green